OpenStudy (anonymous):

Find the coordinates of the other endpoint when you are given the midpoint and one of the endpoints. P1 M P2 (10, 6) (-4, 8) ( , )

OpenStudy (radar):

(-18, 10) would make (-4, 8) the midpoint

OpenStudy (radar):

10 - (-4)=14 -4-14=-18 10-8=2 8+2=10

OpenStudy (anonymous):

can u help me with the next 2 questions i posted up

OpenStudy (radar):

O.K I will go to them.